OpenStudy (dominirican1013):

Which of the following vectors are orthogonal to (2,1)? A. (1,2) B. (-3,6) C. (-2,-3) D. (1,-2)

OpenStudy (dominirican1013):

@ganeshie8

OpenStudy (dominirican1013):

@jamesr

OpenStudy (dominirican1013):

@dan815

OpenStudy (anonymous):

YOLO

OpenStudy (dominirican1013):

-.-

OpenStudy (anonymous):

yess

OpenStudy (dominirican1013):

I need help

OpenStudy (anonymous):

ight whts up

OpenStudy (dominirican1013):

Is it B and D then

OpenStudy (chillout):

Do you remember the condition to be orthogonal?

OpenStudy (chillout):

I.E. the dot product.

OpenStudy (anonymous):

aww ship i got to catch up

OpenStudy (dominirican1013):

Yes but is B and D

OpenStudy (anonymous):

i say its D

OpenStudy (dominirican1013):

It was right thanks though

OpenStudy (anonymous):

what, D

OpenStudy (chillout):

\[\vec{a}^{\,}\cdot\vec{b}^{\,}=0 \rightarrow \sum_{1}^{i}a_{i}b_{i} = 0\] This is what you need to use.

OpenStudy (dan815):

hey orthogonal means these vectors are at 90 degree angle

OpenStudy (dan815):

othogonal or perpendicular

OpenStudy (dan815):

the slope for the vector given <2,1> Slope= 1/2 = and a perpendicular slope to that would be 2/-1 =-2 which is the "negative reciprocal" of 1/2 we can see the slopes of each vectors given see which of those simplify to -2 slope A)(1,2) slope = 2/1=2 B. (-3,6) Slope=6/-3=2/-1=-2 <-------- C. (-2,-3) Slope= -3/-2=3/2 D. (1,-2) Slope=-2/1 =-2 <------------ So you are right